# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 6.3. Unramified computation (case2 & case3)
$\left[\begin{smallmatrix}\varpi^{n}a1_{2}&0\\ Y_{\eta}^{\dagger}&\varpi^{e}1_{2}\end{smallmatrix}\right]=\mathsf{n}(\varpi^{n}a(\det Y_{\eta})^{-1}Y_{\eta})\left[\begin{smallmatrix}\tau&0\\ 0&{}^{t}\tau^{\dagger}\end{smallmatrix}\right]h(e+n-2j,j)k$
where $k=\left[\begin{smallmatrix}A&0\\ 0&{}^{t}A^{\dagger}\end{smallmatrix}\right]\left[\begin{smallmatrix}a\varpi^{j}(\det Y_{\eta})^{-1}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\left[\begin{smallmatrix}0&-1_{2}\\ 1_{2}&\varpi^{e}(\det Y_{\eta})^{-1}Y_{\eta}\end{smallmatrix}\right]\in K$. Since $\Lambda$ is unramified and $B$ is $K$-invariant, we have
$=\psi(\varpi^{n}a(\det Y_{\eta})^{-1}\cdot\tfrac{{\mathsf{a}}^{2}{\mathsf{d}}}{2})\cdot B\left(h(e+n-2j,j)\right)$
$=\psi\left(-\varpi^{n}a\cdot\tfrac{{\mathsf{a}}^{4}{\mathsf{d}}}{2}\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)^{-1}\right)\cdot B\left(h(e+n-2j,j)\right).\qed$
@@
Suppose that $\pi$ is of type I or IIb. Let $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ be characters such that $\text{cond}(\mu)=e>0$ and $\Lambda$ is unramified. Suppose that $S=\left[\begin{smallmatrix}{\mathsf{a}}&\mathsf{b}/2\\ \mathsf{b}/2&\mathsf{c}\end{smallmatrix}\right]\in{\rm{Sym^{2}}}(F)\cap{\bf GL}_{2}(\mathfrak{o})$ satisfies that $L/F$ is an unramified field extension and $2d\in\mathfrak{o}^{\times}$. When $\phi=1_{{\mathfrak{p}}^{e}\mathfrak{o}_{L}\oplus 1+{\mathfrak{p}}^{e}\mathfrak{o}_{L}}$ and ${\operatorname{Re}}(s)>1$, we have
@@
Let $K^{\#}({\mathfrak{p}}^{e})=\{k^{\#}\in K^{\#}\ |\ k^{\#}\equiv 1_{2}\ ({\rm mod}\ {\mathfrak{p}}^{e}\mathfrak{o}_{L})\}$ be the principal congruence subgroup of level ${\mathfrak{p}}^{e}\mathfrak{o}_{L}$. If $\phi\in\mathcal{S}(L^{2})$ is right $K^{\#}({\mathfrak{p}}^{e})$-invariant, the function
@@
$=[K^{\#}:K_{0}^{\#}({\mathfrak{p}}^{e})]^{-1}\sum_{[\gamma]\in K_{0}^{\#}({\mathfrak{p}}^{e})\backslash K^{\#}}f_{\phi}^{(s,\Lambda,\mu)}(\gamma)\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\gamma)b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
(6.29) $=q^{-2e+2}(q^{2}+1)^{-1}\sum_{[\gamma]\in K_{0}^{\#}({\mathfrak{p}}^{e})\backslash K^{\#}}f_{\phi}^{(s,\Lambda,\mu)}(\gamma)\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\gamma)b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a.$
@@
(6.30) $\left[\begin{smallmatrix}1&0\\ \xi&1\end{smallmatrix}\right]\ (\xi\in{\mathfrak{p}}\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}),\quad\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right]\ (\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}).$
Now we prove the first equation. By Lemma 6.3, when $\gamma$ runs through the above elements, then we have $f_{\phi}^{(s,\Lambda,\mu)}(1_{2})=q^{-2e+2}(q^{2}-1)^{-1}$ and $f_{\phi}^{(s,\Lambda,\mu)}(\gamma)=0$ if $\gamma\neq 1$. By substituting this in (6.29) and Lemma 6.1, we have
@@
$=q^{-4e+4}(q^{4}-1)^{-1}\int_{F^{\times}}B_{\pi}^{0}\left(\left[\begin{smallmatrix}a\varpi^{e}1_{2}&aS^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]\right)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=q^{-4e+4}(q^{4}-1)^{-1}\int_{F^{\times}}B_{\pi}^{0}\left(\left[\begin{smallmatrix}a\varpi^{e}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\right)\psi(a{\operatorname{tr}}(SS^{\dagger}))\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
@@
(6.31) ×$\times\sum_{\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}}\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right])b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a.$
@@
×$\times W_{L}(\mu_{L},\psi_{L})W_{F}(\mu,\psi)\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}^{e}\mathfrak{o}_{L}\\ {\mathsf{a}}^{6}{\mathsf{d}}+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)$
@@
Let $\pi$, $S$, $\mu$, and $\Lambda$ be as in Proposition 6.6, then
@@
6.4. Computation for old forms of type ${\rm I}$ and ${\rm IIb}$ (case4)
Now we assume that $\pi$ has the trivial character. Then $\pi$ has the $(S,\Lambda)$-Bessel model if and only if $\Lambda=1$. Recall that the dimension of $V_{\pi}^{K_{0}({\mathfrak{p}})}$ is equal to $4$ or $3$ according as $\pi$ is of type I or type IIb.
@@ 6.4. Computation for old forms of type ${\rm I}$ and ${\rm IIb}$ (case4)
We will give the proof of the case that $\pi$ is of type I. In the case of type IIb, the proof is similar.
Let $\{B_{j}^{*}\}_{1{\leqslant}j{\leqslant}4}\subset{\mathscr{B}}(S,1)[\pi]^{K_{0}({\mathfrak{p}})}$ be an eigenbasis of $T_{1,0}$ with eigenvalues $\{\mu_{j}\}_{1{\leqslant}i{\leqslant}4}$ and $(\eta_{ij})_{1{\leqslant}i,j{\leqslant}4}$ be the representation matrix of $\eta|_{K_{0}({\mathfrak{p}})}$ with respect to $\{B_{j}^{*}\}_{1{\leqslant}i{\leqslant}4}$. Then, the first relation in (6.40) shows that
$B_{j}^{*}(h(l+1,0))=\mu_{j}q^{-3}B_{j}^{*}(h(l,0)),\quad l\in{\mathbb{Z}}_{{\geqslant}0}.$
It is easy to check that
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})=L(s+1,\Lambda\mu_{L}),\quad k^{\#}\in K^{\#}.$
By the left $K_{0}^{\#}({\mathfrak{p}})$-coset decomposition of $K^{\#}$;
(6.32) $K^{\#}=K_{0}^{\#}({\mathfrak{p}})\sqcup\left(\sqcup_{\xi\in\mathfrak{o}_{L}/{\mathfrak{p}}\mathfrak{o}_{L}}\left[\begin{smallmatrix}1&\xi\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]K_{0}^{\#}({\mathfrak{p}}^{e})\right),$
the zeta integral (6.8) becomes
(6.33) $Z(\phi,B,s,\mu;\eta)$ $=Z(\phi,\eta B,s,\mu;1_{4})$
$=[K^{\#}:K_{0}^{\#}({\mathfrak{p}})]^{-1}L(s+1,\Lambda\mu_{L})$
×$\times\int_{F^{\times}}\left\{\eta B(\mathsf{m}(a1_{2},a))+\sum_{\xi\in\mathfrak{o}_{L}/{\mathfrak{p}}\mathfrak{o}_{L}}\eta B\left(\mathsf{m}(a1_{2},a)\iota\left(\left[\begin{smallmatrix}1&\xi\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]\right)\right)\right\}\mu^{-1}(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=\frac{1}{q^{2}+1}L(s+1,\Lambda\mu_{L})$
×$\times\left\{\sum_{l\in{\mathbb{Z}}}\eta B(h(l,0))+q^{2}\Lambda(\varpi)^{-1}B(h(l+1,0))\right\}\mu(\varpi)^{l}q^{-l(s-1)}.$
Since $B(h(l,0))=0$ if $l<0$ by (6.25), the right hand side equals
(6.34) $\frac{L(s+1,\Lambda\mu_{L})}{q^{2}+1}\left\{\sum_{l=0}^{\infty}\eta B(h(l,0))\mu(\varpi)^{l}q^{-l(s-1)}+q^{s+1}\Lambda(\varpi)^{-1}\mu(\varpi)^{-1}\sum_{l=0}^{\infty}B(h(l,0))\mu(\varpi)^{l}q^{-l(s-1)}\right\}.$
By substituting $B=B_{j}^{*}$ in (6.34) and $\Lambda=1$,
$Z(\phi,B_{j}^{*},s,\mu;\eta)=\frac{L(s+1,\Lambda\mu_{L})}{q^{2}+1}\left\{\sum_{i=1}^{4}\frac{\eta_{ij}B_{i}^{*}(1_{4})}{1-\mu_{j}\mu(\varpi)q^{-(s+2)}}+\frac{\mu(\varpi)^{-1}q^{s+1}B_{j}^{*}(1_{4})}{1-\mu_{i}\mu(\varpi)q^{-(s+2)}}\right\}.$
@@
$Z(\phi,B_{j}^{*},s,\mu;\eta)$ $=\frac{1}{q^{2}+1}\prod_{j=1}^{4}(1-\mu_{j}q^{-2}X)^{-1}$
×$\times\frac{A_{0}+A_{1}X+A_{2}X^{2}+A_{3}X^{3}+A_{4}X^{4}}{X(1-q^{-2}X^{2})},\quad(X=\mu(\varpi)q^{-s},A_{i}\in{\mathbb{C}},1{\leqslant}i{\leqslant}4).$
By noting that $\{\mu_{1},\mu_{2},\mu_{3},\mu_{4}\}=\{\alpha\beta\gamma q^{3/2},\beta\gamma q^{3/2},\alpha\gamma q^{3/2},\gamma q^{3/2}\}$ from the table (6.15), we have
(6.35) $(q^{2}+1)\cdot\frac{Z(\phi,B_{j}^{*},s,\mu;1_{4})}{L(s+1/2,\pi,\mu)}=\frac{A_{0}+A_{1}X+A_{2}X^{2}+A_{3}X^{3}+A_{4}X^{4}}{X(1-q^{-2}X^{2})}$
@@
$A_{1}=\sum_{i=1}^{4}\eta_{ij}B_{i}^{*}(1_{4})+qB_{j}^{*}(1_{4})\left\{\left(\sum_{k=1}^{4}q^{-2}\mu_{k}\right)-q^{-2}\mu_{i}\right\}$
(6.37) $=\eta B_{j}^{*}(1_{4})+q^{-1}{\operatorname{tr}}(T_{1,0}|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\cdot B_{j}^{*}(1_{4})-q^{-1}\cdot T_{1,0}B_{j}^{*}(1_{4}).$
The functional equation (6.9) and the fact $\varepsilon(\pi,s,\mu,\psi)=1$ from ([30]) show that the right hand side of (6.35) is invariant under the transformation $X\to X^{-1}$. By comparing the coefficients, we have
(6.38) $A_{2}=q^{-2}(q^{2}-1)A_{0},\quad A_{3}=-q^{-2}A_{1},\quad A_{4}=-q^{-2}A_{0}$
Substituting (6.36), (6.37), and (6.38) into (6.35), we get
$(q^{2}+1)\cdot\frac{Z(\phi,B_{j}^{*},s,\mu;\eta)}{L(s+1/2,\pi,\mu)}$
$=\frac{A_{0}+A_{1}X+q^{-2}(q^{2}-1)A_{0}X^{2}-q^{-2}A_{1}X^{3}-q^{-2}A_{0}X^{4}}{X(1-q^{-2}X^{2})}$
$=A_{0}(X+X^{-1})+A_{1}$
$=\left[\eta B+q^{-1}T_{1,0}B+\left\{\mu(\varpi)^{-1}q^{s+1}+\mu(\varpi)q^{-s+1}-{\operatorname{tr}}(q^{-1}T_{1,0}|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\right\}B\right](1_{4}).\qed$
6.5. Computation for newforms of type ${\rm III}a$ (case5)
By [32, §9.1], there exists a basis $\{B_{1},B_{2}\}$ of ${\mathscr{B}}(S,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ such that

### v2

@@ 6.3. Unramified computation (case 2 and case 3)
$\left[\begin{smallmatrix}\varpi^{n}a1_{2}&0\\ Y_{\eta}^{\dagger}&\varpi^{e}1_{2}\end{smallmatrix}\right]=\mathsf{n}(\varpi^{n}a(\det Y_{\eta})^{-1}Y_{\eta})\left[\begin{smallmatrix}\tau&0\\ 0&{}^{t}\tau^{\dagger}\end{smallmatrix}\right]h(e+n-2j,j)k,$
where $k=\left[\begin{smallmatrix}A&0\\ 0&{}^{t}A^{\dagger}\end{smallmatrix}\right]\left[\begin{smallmatrix}a\varpi^{j}(\det Y_{\eta})^{-1}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\left[\begin{smallmatrix}0&-1_{2}\\ 1_{2}&\varpi^{e}(\det Y_{\eta})^{-1}Y_{\eta}\end{smallmatrix}\right]\in K$. Since $\Lambda$ is unramified and $B$ is $K$-invariant, we have the desired equality. ∎
@@
Suppose that $\pi$ is of type I or IIb. Let $\mu:F^{\times}\to{\mathbb{C}}^{1}$ and $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$ be characters such that $\text{cond}(\mu)=e>0$ and $\Lambda$ is unramified. Suppose that $T=\left[\begin{smallmatrix}{\mathsf{a}}&\mathsf{b}/2\\ \mathsf{b}/2&\mathsf{c}\end{smallmatrix}\right]\in{\rm{Sym^{2}}}(F)\cap{\bf GL}_{2}(\mathfrak{o})$ satisfies that $L/F$ is an unramified field extension and $2d\in\mathfrak{o}^{\times}$. When $\phi=1_{{\mathfrak{p}}_{L}^{e}\oplus 1+{\mathfrak{p}}_{L}^{e}}$ and ${\operatorname{Re}}(s)>1$, we have
@@
Let $K^{\#}({\mathfrak{p}}^{e})=\{k^{\#}\in K^{\#}\ |\ k^{\#}\equiv 1_{2}\ ({\rm mod}\ {\mathfrak{p}}_{L}^{e})\}$ be the principal congruence subgroup of level ${\mathfrak{p}}_{L}^{e}$. If $\phi\in\mathcal{S}(L^{2})$ is right $K^{\#}({\mathfrak{p}}^{e})$-invariant, the function
@@
(6.24) $=[K^{\#}:K_{0}^{\#}({\mathfrak{p}}^{e})]^{-1}\sum_{[\gamma]\in K_{0}^{\#}({\mathfrak{p}}^{e})\backslash K^{\#}}f_{\phi}^{(s,\Lambda,\mu)}(\gamma)\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\gamma)b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a.$
@@
(6.25) $\left[\begin{smallmatrix}1&0\\ \xi&1\end{smallmatrix}\right]\ (\xi\in{\mathfrak{p}}_{L}/{\mathfrak{p}}_{L}^{e}),\quad\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right]\ (\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}_{L}^{e}).$
Now we prove the first equation. By Lemma 6.3, when $\gamma$ runs through the above elements, then we have $f_{\phi}^{(s,\Lambda,\mu)}(1_{2})=q^{-2e+2}(q^{2}-1)^{-1}$ and $f_{\phi}^{(s,\Lambda,\mu)}(\gamma)=0$ if $\gamma\neq 1$. By substituting this in (6.24), noting that $[K^{\#}:K_{0}^{\#}({\mathfrak{p}}^{e})]=q^{2e-2}(q^{2}+1)$, and Lemma 6.1, we have
@@
$=q^{-4e+4}(q^{4}-1)^{-1}\int_{F^{\times}}B_{\pi}^{0}\left(\left[\begin{smallmatrix}a\varpi^{e}1_{2}&aT^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]\right)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=q^{-4e+4}(q^{4}-1)^{-1}\int_{F^{\times}}B_{\pi}^{0}\left(\left[\begin{smallmatrix}a\varpi^{e}1_{2}&0\\ 0&1_{2}\end{smallmatrix}\right]\right)\psi(a\cdot{\operatorname{tr}}(TT^{\dagger}))\mu(a)|a|^{s-1}{\rm{d}}^{\times}a$
@@
(6.26) ×$\times\sum_{\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}_{L}^{e}}\int_{F^{\times}}B_{\pi}^{0}(\mathsf{m}(a1_{2},a)\iota(\left[\begin{smallmatrix}0&-1\\ 1&\eta^{\dagger}\end{smallmatrix}\right])b)\mu(a)|a|^{s-1}{\rm{d}}^{\times}a.$
@@
×$\times W_{L}(\mu_{L},\psi_{L})W_{F}(\mu,\psi)\sum_{\begin{subarray}{c}\eta\in\mathfrak{o}_{L}/{\mathfrak{p}}_{L}^{e}\\ {\mathsf{a}}^{6}{\mathsf{d}}+{\operatorname{N}}_{L/F}(\eta)\in\mathfrak{o}^{\times}\end{subarray}}\mu\left(\tfrac{{\mathsf{a}}^{6}{\mathsf{d}}}{4}+{\operatorname{N}}_{L/F}(\eta)\right)$
@@
Let $\pi$, $T$, $\mu$, and $\Lambda$ be as in Proposition 6.6, then
@@
6.4. Unramified computation (case 4 and case 5)
Suppose that $f>0$ and $L/F$ is inert. The section $f_{\phi}^{(s,\Lambda,\mu)}$ restricted to $K^{\#}$ is explicitly determined as
$f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})=\begin{cases}q^{-2f+2}(q^{2}-1)^{-1}\Lambda(d)\quad&(c\in{\mathfrak{p}}_{L}^{e}),\\ 0&({\rm otherwise}),\end{cases}\quad k^{\#}=\left[\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right]\in K^{\#}$
by a direct calculation.
Proposition 6.8.
Suppose that $\pi$ is of type I or IIb. Let $\mu$, $\Lambda$, $T$ be as above. When $\phi=\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{{\mathfrak{p}}_{L}^{f}\oplus 1+{\mathfrak{p}}_{L}^{f}}$ and ${\operatorname{Re}}(s)>1$, we have
$Z(\phi,B_{\pi}^{0},s,\mu;b^{\prime})=Z(\widehat{\phi},B_{\pi}^{0},s,\mu;b^{\prime})=q^{-4f+4}(q^{4}-1)^{-1}L(s+1/2,\pi,\mu).$
Proof.
It suffices to calculate the integral $Z(\phi,B_{\pi}^{0},s,\mu;b^{\prime})$ due to (6.9) and the fact that $\varepsilon(\pi,s,\mu,\psi)=1$ ([31, Theorem 4.4]). For any $k^{\#}=\left[\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right]\in K_{0}^{\#}({\mathfrak{p}}^{f})$, it can be checked that $k^{\#}=\left[\begin{smallmatrix}\tau&\\ &\lambda\tau^{\dagger}\end{smallmatrix}\right]\left[\begin{smallmatrix}1&\beta\\ &1\end{smallmatrix}\right]\left[\begin{smallmatrix}1&\\ cd^{-1}&1\end{smallmatrix}\right]$, where we set $(\tau,\lambda,\beta)\in{\mathfrak{o}}_{L}^{\times}\times{\mathfrak{o}}^{\times}\times{\mathfrak{o}}_{L}$ as
$\tau=(\det k^{\#})^{-1}d,\quad\lambda=(\det k^{\#})(dd^{\dagger}),\quad\beta=-b/d.$
This and a short calculation yield the proposition due to [44, Theorem 2.1]. ∎
6.5. Computation for old forms of type ${\rm I}$ and ${\rm IIb}$ (case 6)
Now we assume that $\pi$ has the trivial character. Then, $\pi$ has the $(T,\Lambda)$-Bessel model if and only if $\Lambda=1$. Recall that the dimension of $V_{\pi}^{K_{0}({\mathfrak{p}})}$ is equal to $4$ or $3$ according as $\pi$ is of type I or type IIb.
@@ 6.5. Computation for old forms of type ${\rm I}$ and ${\rm IIb}$ (case 6)
We will give a proof of the case that $\pi$ is of type I. In the case of type IIb, the proof is similar. Let $\{B_{j}^{*}\}_{1{\leqslant}j{\leqslant}4}\subset{\mathscr{B}}(T,1)[\pi]^{K_{0}({\mathfrak{p}})}$ be an eigenbasis of $T_{1,0}$ with eigenvalues $\{\lambda_{j}\}_{1{\leqslant}i{\leqslant}4}$. Then, we suffice to check the desired equation only for $B=B_{j}^{*}$ $(1{\leqslant}j{\leqslant}4)$. Let $(\eta_{ij})_{1{\leqslant}i,j{\leqslant}4}$ be the representation matrix of $\eta|_{K_{0}({\mathfrak{p}})}$ with respect to $\{B_{j}^{*}\}_{1{\leqslant}i{\leqslant}4}$. Then, the first relation in (6.34) shows that
$B_{j}^{*}(h(l+1,0))=\lambda_{j}q^{-3}B_{j}^{*}(h(l,0)),\quad l\in{\mathbb{Z}}_{{\geqslant}0}.$
It is easy to check the equality $f_{\phi}^{(s,\Lambda,\mu)}(k^{\#})=L(s+1,\Lambda\mu_{L})$ for all $k^{\#}\in K^{\#}$. By considering the left $K_{0}^{\#}({\mathfrak{p}})$-coset decomposition of $K^{\#}$;
(6.27) $K^{\#}=K_{0}^{\#}({\mathfrak{p}})\sqcup\left(\sqcup_{\xi\in\mathfrak{o}_{L}/{\mathfrak{p}}_{L}}\left[\begin{smallmatrix}1&\xi\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]K_{0}^{\#}({\mathfrak{p}}^{e})\right),$
the zeta integral $Z(\phi,B,s,\mu;\eta)$ becomes $[K^{\#}:K_{0}^{\#}({\mathfrak{p}})]^{-1}L(s+1,\Lambda\mu_{L})$ times
$\int_{F^{\times}}\left\{\eta B(\mathsf{m}(a1_{2},a))+\sum_{\xi\in\mathfrak{o}_{L}/{\mathfrak{p}}_{L}}\eta B\left(\mathsf{m}(a1_{2},a)\iota\left(\left[\begin{smallmatrix}1&\xi\\ 0&1\end{smallmatrix}\right]\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]\right)\right)\right\}\mu^{-1}(a)|a|^{s-1}{\rm{d}}^{\times}a$
$=\sum_{l\in{\mathbb{Z}}}\left\{\eta B(h(l,0))+q^{2}B(h(l+1,0))\right\}\mu(\varpi)^{l}q^{-l(s-1)}$
(6.28) $=\sum_{l=0}^{\infty}\eta B(h(l,0))\mu(\varpi)^{l}q^{-l(s-1)}+q^{s+1}\mu(\varpi)^{-1}\sum_{l=0}^{\infty}B(h(l,0))\mu(\varpi)^{l}q^{-l(s-1)}$
because $B(h(l,0))=0$ if $l<0$ by (6.21). By substituting $B=B_{j}^{*}$ in (6.28) and $\Lambda=1$, we have
$Z(\phi,B_{j}^{*},s,\mu;\eta)=\frac{L(s+1,\Lambda\mu_{L})}{q^{2}+1}\left\{\sum_{i=1}^{4}\frac{\eta_{ij}B_{i}^{*}(1_{4})}{1-\lambda_{j}\mu(\varpi)q^{-(s+2)}}+\frac{\mu(\varpi)^{-1}q^{s+1}B_{j}^{*}(1_{4})}{1-\lambda_{i}\mu(\varpi)q^{-(s+2)}}\right\}.$
@@
$Z(\phi,B_{j}^{*},s,\mu;\eta)$ $=\frac{1}{q^{2}+1}\prod_{j=1}^{4}(1-\lambda_{j}q^{-2}X)^{-1}\frac{A_{0}+A_{1}X+A_{2}X^{2}+A_{3}X^{3}+A_{4}X^{4}}{X(1-q^{-2}X^{2})}.$
with $X=\mu(\varpi)q^{-s}$ for some $A_{i}\in{\mathbb{C}}\,(1{\leqslant}i{\leqslant}4)$.
By noting that $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}=\{\alpha\beta\gamma q^{3/2},\beta\gamma q^{3/2},\alpha\gamma q^{3/2},\gamma q^{3/2}\}$ from [36, Table A.8], we have
(6.29) $(q^{2}+1)\cdot\frac{Z(\phi,B_{j}^{*},s,\mu;1_{4})}{L(s+1/2,\pi,\mu)}=\frac{A_{0}+A_{1}X+A_{2}X^{2}+A_{3}X^{3}+A_{4}X^{4}}{X(1-q^{-2}X^{2})}.$
@@
(6.31) $A_{1}=\eta B_{j}^{*}(1_{4})+q^{-1}{\operatorname{tr}}(T_{1,0}|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\cdot B_{j}^{*}(1_{4})-q^{-1}\cdot T_{1,0}B_{j}^{*}(1_{4}).$
The functional equation (6.9) and the relation $\varepsilon(\pi,s,\mu,\psi)=1$ ([31, Theorem 4.4]) show that the right-hand side of (6.29) is invariant under the transformation $X\to X^{-1}$. By comparing the coefficients, we have
(6.32) $A_{2}=q^{-2}(q^{2}-1)A_{0},\quad A_{3}=-q^{-2}A_{1},\quad A_{4}=-q^{-2}A_{0}.$
Substituting (6.30), (6.31), and (6.32) into (6.29), the right-hand side of (6.29) equals
$\left[\eta B_{j}^{*}+q^{-1}T_{1,0}B_{j}^{*}+\left\{\mu(\varpi)^{-1}q^{s+1}+\mu(\varpi)q^{-s+1}-{\operatorname{tr}}(q^{-1}T_{1,0}|_{V_{\pi}^{K_{0}({\mathfrak{p}})}})\right\}B_{j}^{*}\right](1_{4}).$
6.6. Computation for newforms of type ${\rm III}a$ (case 7)
By [33, §9.1], there exists a basis $\{B_{1},B_{2}\}$ of ${\mathscr{B}}(T,\Lambda)[\pi]^{K_{0}({\mathfrak{p}})}$ such that